addie0506:

the population of a species is modeled by the equation p(t) = -t^4+72t^2+225, where t is the number of years. find the approximate number of years until the species is extinct

itsme:

immm dumb at maths @supie

supie:

This post is closed we don't have to answer it. Anyways ima jus do it anyways So what it said in the equation was \[p(t) = -t^4+72t^2+225\] but we are going to have to do \[p(t) = -t^4+72t^2+225=0\] Since \[P(t)=0\] So we solve for t so \[(t^2+3)(t^2-75)=0\]\[t^4−72t^2−225=0\]so we factor, specifically the left side of the equation: \[(t^2+3)(t^2−75)=0\] \[t^2+3=0\]=\[t^2−75=0\] so now that should give us the answer of \[t=?\]